Define a function A: N x N -> N as follows: A(m, n) ={2n, if m= 0; 0, if m≥1 and n= 0; 2, if m≥1 and n= 1; A(m-1, A(m, n-1)), if m≥1 and n≥2

Show that A(1, n) = 2^n whenever n≥1.
Hint: Induct on n.
1
Expert's answer
2020-12-15T02:19:03-0500

We shall prove by inducting over "n".

When "n=1" .

"A(1,1)=2=2^1."

We see that it is true for "n=1".


Suppose it is true for some "n=k\geq1" . That is, "A(1,k)=2^k" .

We shall prove that it holds for "n=k+1" . Since "k\geq1 \implies k+1\geq2" .

So,

"A(1,k+1)=A(1-1,A(1,k+1-1))\\
A(1,k+1)=A(0,A(1,k))\\"

Recall that "A(1,k)=2^k" .


"\implies A(1,k+1)=A(0,2^k)=2.2^k=2^{k+1}" .


Therefore, "A(1,k+1)=2^{k+1}" .


This shows that it is true for "n=k+1" . Hence, "A(1,n)=2^n" whenever "n\geq1" .